1. This morning, Fish Crown Inc.decided to take some leverage for the very firsttime in the company’s history.Specifically, the company just issued 3-year zero-coupon bonds with a face value of $1,000. All bonds were priced at $800.Following the issue of the bonds, the company’sWACC is calculated at 10%.The company’s tax rate is 35% and therequired return on equity is15%.What fractionof the Fish Crown Inc’sassetsisinnet debt, and what fractionis in equity?

2. Anheuser-Bush InBevsustains in the last years a fixed level of debt of D=$110B, which is regarded to be risk-free. Moreover, the market value of its shareholders’ equity is E=$70B and the beta of its stock βE=1.4.Suppose that the CAPM holds. The risk-free rate is 2% and the market risk premium is 5%.The company plans to introduce a new ultra-light beer with zero calories, which will be called BudZero. The cost of bringing the beer to the market (at t=0) is $200M.BudZerois then expected to generate constantfree cash flows equal to $100M, paid at the end of every year (starting at t=1),forever. Suppose that the “BudZero” project will be financed exclusively with equity.What is the net present value of this project?

3. Nike, Inc. is considering a $10M investment (paid at t=0) in new “smarter” sneakers.Thenew sneakers areexpected to generate free cash flows of $3M, at the end of every year(starting at t=1), for the next6 years.They will be totally out of fashion after that.Nike will finance this project 20% with debt and 80% with equity.Assume a risk-free rate of 2% anda market-risk premium of 5%.The corporation’s equity beta in the stock market is 1.2. Its stock price is $60 per share and, overall, there are 2Bshares outstanding. The corporation has a current debt of $13.5Band holds $1.5Bin cash.The corporation’s debt is risk-freewith a yield-to-maturity of 2%. The corporate tax rate is 40%. What is the net present value of the project?

4. Consider a portfolio that consistsof shares of Ford and T-Bills.Ford’s stockhas an expected return μFord= 2% and a standard deviation σFord=6.5%.T-Bills are considered risk-freeand offer a surereturn of 1%.Suppose that the standard deviation of the portfolio is σportfolio=4%.What is the portfolio’s expected return (μportfolio)?
Answer #1

1]

Before tax cost of debt = YTM of bonds

Issue price of bonds = face value / (1 + YTM)years to maturity

$800 = $1,000 / (1 + YTM)3

YTM = (1000 / 800)1/3 - 1

YTM = 7.72%

After tax cost of debt = Before tax cost of debt * (1 - tax rate)

After tax cost of debt = 7.72% * (1 - 35%) = 5.02%

Let us say the fraction of debt is X. Then, the fraction of equity = 1 - X

WACC = (weight of debt * after tax cost of debt) + (weight of common stock * cost of common stock)

10% = (X * 5.02%) + ((1 - X) * 15%)

0.10 = 0.0502X + 0.15 - 0.15X

0.0998X = 0.05

X = 0.50

Fraction of debt = 0.50

Fraction of equity = 1 - 0.50 = 0.50